# Critical behaviour of the fully packed loop-$O(n)$ model on planar triangulations

### v2

@@ 3 Hamburger-cheeseburger and the partition function
$\gamma_{+}=\frac{1}{2x_{c}},$
This extra input will allow us to solve the resolvent equation (2.20) in Section 4. This will go through establishing an exact relation between the partition function $F_{\ell}$ and the marginal law of the filled-in cluster at the origin in the infinite FK map. This part is close to some of the arguments in [da2025scaling].
@@ 3.1 Typical clusters as skeleton words
We now show that the reduced walks of Section 2.2.2 carry some geometric information about the FK planar maps. We will be interested in filled-in clusters since these will be easy to relate to the partition function $F_{\ell}$. However, we stress that similar techniques can be leveraged to access further geometric information, such as loop exponents, as done in [berestycki2017critical]. Analogous results for the case $q=4$ (which is not covered here) are presented in [da2025scaling].
@@ 3.3 Dictionary and proof of ansatz
We then use this identity to deduce the value of the right endpoint of the cut $\gamma_{+}$, thereby proving the ansatz underlying the analysis in [BBG12] and discussed in the introduction and preliminaries.
@@
Let $\varphi(z):=W(1/z)=\sum_{\ell\geq 0}F_{\ell}z^{\ell+1}$. By Section 2.3.2, The function $\varphi$ is a power series, let $R_{\varphi}$ be its radius of convergence. Since the coefficients of $\varphi$ are positive we deduce from the Vivanti–Pringsheim theorem that it cannot be continued analytically to any neighbourhood of $R_{\varphi}$. Furthermore, the arguments of [BBG12b] show that it can be continued analytically to $\mathbb{C}\setminus((-\infty,1/\gamma_{-}])\cup[1/\gamma_{+},\infty))$ by definition of $\gamma_{-}$ and $\gamma_{+}$. This implies $R_{\varphi}=1/\gamma_{+}$ necessarily (and $|\gamma_{-}|\leq\gamma_{+}$ as already announced).
@@
Using [berestycki2017critical, Lemma A.2] (see also the end of the proof of Theorem 1.1 there), one can see that
@@
for some constant $C>0$. Actually, we will deduce more precise asymptotics and even an exact expression for $F_{\ell}$, using a completely different method (and without using [berestycki2017critical]): namely, we will solve the resolvent equation (2.20) explicitly, relying on 3.6. This is the purpose of the next section. Note that translating back through 3.5, this will also imply more precise asymptotics than in [berestycki2017critical], and even an exact expression of, say, the marginals of $\tau^{\mathsf{h}}$. This will be carried out in Section 5.
@@ 4 Solution to the resolvent equation in the self-dual case
Importantly, remark that the set of equations (E)–(N) is the same as that of Gaudin and Kostov [gaudin1989n] (see equations (10) and (11) there, with $g=0$). We now solve this equation for fixed $\gamma_{-},\gamma_{+}$ making mathematical sense of the ideas of [gaudin1989n]. Recall from 3.6 that we identified $x_{c}=\frac{1}{2\gamma_{+}}$.
@@
For the second item, we use the expression of $k$ in (4.2), from which we see that $k$ is continuous except at 0 where it has a $1/z$ singularity, and $k(z)=O(\mathrm{e}^{-|z|})$ as $|z|\to\infty$. Thus, $K$ is indeed well-defined as a Cauchy principal value on $\{z\in\mathbb{C},-1<\Im(z)<1\}$. Moreover, it is known that $K$ has the said explicit expression, see e.g. formulas F60-F61 in [raade1995mathematics, Chapter 13] (and so, in particular, it is analytic).
@@
with [^1] Here we stress that [gaudin1989n, Equation (32)] has a typo: the term $\frac{1-i\omega}{4}$ in the Gamma function should be replaced by $\frac{1-i\omega}{2}$.
@@ Conclusion.
We will prove this identity in Appendix A (note that this differs from the analogous identity in [gaudin1989n, (38)], though their final answer appears to be correct). Undoing the change of variables, we conclude that for $y\in(\gamma_{-},\gamma_{+})$,
@@
5 Loop and cluster exponents
@@ Appendix A Computation of inverse Fourier transform
Such integrals can be explicitly computed in terms of Beta functions: indeed, if $a,b\in\mathbb{C}$ with $\Re(a)>|\Re(b)|$ then (see [DLMF, (5.12.7)])

### v3

@@ 3 Hamburger-cheeseburger and the partition function
$\gamma_{+}=\frac{1}{2x_{c}}.$
This extra input will allow us to solve the (colour-symmetric) resolvent equation (2.23) in Section 4. This will go through establishing an exact relation between the partition function $F_{\ell}$ and the marginal law of the filled-in cluster at the origin in the infinite FK map. This part is close to some of the arguments in [DHP+25].
@@ 3.1 Typical clusters as skeleton words
We now show that the reduced walks of Section 2.2.2 carry some geometric information about the FK planar maps. We will be interested in filled-in clusters since these will be easy to relate to the partition function $F_{\ell}$. However, we stress that similar techniques can be leveraged to access further geometric information, such as loop exponents, as done in [BLR17]. Analogous results for the case $q=4$ (which is not covered here) are presented in [DHP+25].
@@ 3.3 Dictionary and proof of ansatz
We then use this identity to deduce the value of the right endpoint of the cut $\gamma_{+}$, thereby proving the ansatz underlying the analysis in [BBG12b] and discussed in the introduction and preliminaries.
@@
Let $\varphi(z):=W(1/z)=\sum_{\ell\geq 0}F_{\ell}z^{\ell+1}$. By Section 2.3.2, The function $\varphi$ is a power series, let $R_{\varphi}$ be its radius of convergence. Since the coefficients of $\varphi$ are positive we deduce from the Vivanti–Pringsheim theorem that it cannot be continued analytically to any neighbourhood of $R_{\varphi}$. Furthermore, the arguments of [BBG12c] show that it can be continued analytically to $\mathbb{C}\setminus((-\infty,1/\gamma_{-}])\cup[1/\gamma_{+},\infty))$ by definition of $\gamma_{-}$ and $\gamma_{+}$. This implies $R_{\varphi}=1/\gamma_{+}$ necessarily (and $|\gamma_{-}|\leq\gamma_{+}$ as already announced).
@@
Using [BLR17, Lemma A.2] (see also the end of the proof of Theorem 1.1 there), one can see that
@@
for some constant $C>0$. Actually, we will deduce more precise asymptotics and even an exact expression for $F_{\ell}$, using a completely different method (and without using [BLR17]): namely, we will solve the resolvent equation (2.23) explicitly, relying on 3.6. This is the purpose of the next section. Note that translating back through 3.5, this will also imply more precise asymptotics than in [BLR17], and even an exact expression of, say, the marginals of $\tau^{\mathsf{h}}$. This will be carried out in Section 5.
@@ 4 Solution to the resolvent equation in the self-dual case
Importantly, remark that the set of equations (E)–(N) is the same as that of Gaudin and Kostov [GK89] (see equations (10) and (11) there, with $g=0$). We now solve this equation for fixed $\gamma_{-},\gamma_{+}$ making mathematical sense of the ideas of [GK89]. Recall from 3.6 that we identified $x_{c}=\frac{1}{2\gamma_{+}}$.
@@
For the second item, we use the expression of $k$ in (4.2), from which we see that $k$ is continuous except at 0 where it has a $1/z$ singularity, and $k(z)=O(\mathrm{e}^{-|z|})$ as $|z|\to\infty$. Thus, $K$ is indeed well-defined as a Cauchy principal value on $\{z\in\mathbb{C},-1<\Im(z)<1\}$. Moreover, it is known that $K$ has the said explicit expression, see e.g. formulas F60-F61 in [RW95, Chapter 13] (and so, in particular, it is analytic).
@@
with [^1] Here we stress that [GK89, Equation (32)] has a typo: the term $\frac{1-i\omega}{4}$ in the Gamma function should be replaced by $\frac{1-i\omega}{2}$.
@@ Conclusion.
We will prove this identity in Appendix A (note that this differs from the analogous identity in [GK89, (38)], though their final answer appears to be correct). Undoing the change of variables, we conclude that for $y\in(\gamma_{-},\gamma_{+})$,
@@
5 Exact loop and cluster exponents at the self-dual point
@@
6 Exponential decay away from the self-dual point
In this section, we mainly focus on the FK($q$) planar map model away from self-duality and prove 1.4. We take the parameters $(s,p_{0})$ in the region $\mathcal{S}_{q}$ defined in (1.9). Recall that, through the correspondence described in Section 2.3.1, the FK($q$) model (1.1) with percolation parameter $p_{0}\neq p_{\mathrm{sd}}(q)$ is equivalent to a bicoloured fully packed loop-$O(n)$ model on triangulations with non matching face weights $x_{1}$ and $x_{2}$. Choosing $(s,p_{0})\in\mathcal{S}_{q}$ then corresponds to $x_{1},x_{2}\leq x_{c}$ and $(x_{1},x_{2})\neq(x_{c},x_{c})$, as noted at the end of Section 2.3.1. Henceforth, we thus consider the bicoloured loop-$O(n)$ model with weights $x_{1}\leq x_{2}$ satisfying these relations (note that we can always assume $x_{1}\leq x_{2}$ by swapping the roles of colours $1$ and $2$). From now on, the weights $x_{1}$ and $x_{2}$ are fixed as above.
Let us first recall some notation. The notation $\mathfrak{K}$ refers to the cluster of the root triangle (that corresponds to the cluster of the root edge in the original FK model), which may be of either colour. Recall that $K_{\ell}^{(i)}$, $i=1,2$, is the partition function of Fortuin–Kasteleyn maps such that $\mathfrak{K}$ has colour $i$ and perimeter $\ell$. Alternatively — by the above correspondence — this is the partition function of (bicoloured) loop-decorated maps in the loop-$O(n)$ model on triangulations. The purpose of this section is to provide an argument to show that under the above condition on the weights $x_{1}$ and $x_{2}$, the partition function $K_{\ell}^{(i)}$ is exponentially decaying as $\ell\to\infty$, for both $i=1,2$, which is the gist of 1.4. For $i=1,2$ we let $F_{\ell}^{(i)}$ and $\gamma_{+}^{(i)}$ the partition function and right endpoint of the cut corresponding to the colour $i$ in the bicoloured loop-$O(n)$ model with weights $x_{1}$ and $x_{2}$, see our notation in Section 2.3.2.
We first establish some preliminary bounds on $\gamma_{+}^{(i)}$, $i=1,2$. These bounds improve on (2.22).
Lemma 6.1 (Bounds on $\gamma_{+}^{(i)}$ off self-duality).
Assuming $x_{1}<x_{2}\leq x_{c}$, we have the following bounds
$x_{1}\gamma_{+}^{(1)}<\frac{1}{2}\quad\text{and}\quad x_{2}\gamma_{+}^{(2)}\leq\frac{1}{2}.$ (6.1)
If $x_{1}=x_{2}<x_{c}$, we have
$x_{1}\gamma_{+}^{(1)}=x_{2}\gamma_{+}^{(2)}<\frac{1}{2}.$ (6.2)
In any case, for $x_{1}\leq x_{2}\leq x_{c}$ and $(x_{1},x_{2})\neq(x_{c},x_{c})$, we always have
$x_{1}\gamma_{+}^{(1)}+x_{2}\gamma_{+}^{(2)}<1.$
Proof.
We first prove our claim (6.1) on colour $1$. Indeed, using the inequality $x_{1}<x_{2}$ in (2.22), we first deduce that $x_{1}(\gamma_{+}^{(1)}+\gamma_{+}^{(2)})<1$. Let
$\varphi^{(1)}(z)=\sum_{\ell\geq 0}F^{(1)}_{\ell}z^{\ell+1}\quad\text{and}\quad\varphi^{(2)}(z)=\sum_{\ell\geq 0}F^{(2)}_{\ell}z^{\ell+1}.$ (6.3)
By the Vivanti-Pringsheim theorem, the radii of convergence $R^{(1)}$ and $R^{(2)}$ of $\varphi^{(1)}(z)$ and $\varphi^{(2)}(z)$ are $1/\gamma^{(1)}_{+}$ and $1/\gamma^{(2)}_{+}$ respectively. Moreover, since $x_{2}>x_{1}$, $F^{(2)}_{\ell}>F^{(1)}_{\ell}$ and so $R^{(1)}\geq R^{(2)}$, meaning $\gamma^{(1)}_{+}\leq\gamma^{(2)}_{+}$. Hence, from $(\gamma_{+}^{(1)}+\gamma_{+}^{(2)})x_{1}<1$ we conclude that $2\gamma^{(1)}_{+}x_{1}<1$, which proves our claim on colour $1$.
Let us prove our claim (6.1) on colour $2$. Let
$\varphi_{c}(z)=\sum_{\ell\geq 0}F_{\ell}z^{\ell+1},$
where $F_{\ell}$ is the partition function corresponding to the self-dual parameter $x_{c}$. Because $x_{2}\leq x_{c}$, the radius of convergence $R_{c}$ of $\varphi_{c}$ is smaller than $R_{2}$. Again, by the Vivanti-Pringsheim theorem, we deduce the inequality $\gamma_{+}^{(2)}\leq\gamma_{+}$ where $\gamma_{+}$ refers to the self-dual model. In the latter model, we have shown that, in fact, the formula $\gamma_{+}=1/(2x_{c})$ holds (see 3.6). This shows that $\gamma_{+}^{(2)}\leq 1/2x_{c}$. Therefore, $x_{2}\gamma_{+}^{(2)}\leq\tfrac{x_{2}}{2x_{c}}$ and we conclude the proof of our claim on colour $i=2$ by the assumption $x_{2}\leq x_{c}$.
The proof of (6.2) is easier since we clearly have $x_{1}\gamma_{+}^{(1)}=x_{2}\gamma_{+}^{(2)}$ by the condition that $x_{1}=x_{2}$, and then $x_{2}\gamma_{+}^{(2)}<1/2$ repeating the above paragraph. The last claim of 6.1 is a simple consequence of (6.1) and (6.2). ∎
We can now complete the proof of 1.4.
Proof of 1.4.
Let $i=1,2$ and denote by $j$ the complementary colour. By the gasket decomposition, we may express the partition function $K_{\ell}^{(i)}$ as follows:
$K_{\ell}^{(i)}=nF_{\ell}^{(i)}\sum_{\ell^{\prime}\geq 0}A_{\ell,\ell^{\prime}}^{(i\to j)}F_{\ell^{\prime}}^{(j)},$ (6.4)
where $A_{\ell,\ell^{\prime}}^{(i\to j)}$ is the partition function for rings formed of $\ell$ triangles of colour $i$ and $\ell^{\prime}$ triangles of colour $j$ (see Section 2.3.2). We split the proof according to the cases when the parameters are in $\mathcal{S}_{q}$ or in $\mathcal{R}_{q}\setminus\mathcal{S}_{q}$.
Exponential decay in the region $\mathcal{S}_{q}$.
We first prove the part of the statement in 1.4 concerning exponential decay away from self-duality. To do so, we need to bound (6.4) by an exponential term in $\ell$. By definition of the cut, the power series $\varphi^{(1)}$ and $\varphi^{(2)}$ in (6.3) have radii of convergence $1/\gamma_{+}^{(1)}$ and $1/\gamma_{+}^{(2)}$ respectively. This entails that, for any $\varepsilon>0$, the values $\varphi^{(1)}(((1+\varepsilon)\gamma_{+}^{(1)})^{-1})$ and $\varphi^{(2)}(((1+\varepsilon)\gamma_{+}^{(2)})^{-1})$ are finite. In particular, the terms
$\frac{F_{\ell}^{(k)}}{((1+\varepsilon)\gamma_{+}^{(k)})^{\ell}},\quad k=1,2,$
are bounded in $\ell$. Therefore, there exists a constant $C=C(\varepsilon)>0$ such that, for all $\ell\geq 0$,
$F_{\ell}^{(k)}\leq C((1+\varepsilon)\gamma_{+}^{(k)})^{\ell}.$
Let us plug these bounds into (6.4). Then, we get a constant $C=C(\varepsilon)>0$ such that
$K_{\ell}^{(i)}\leq C((1+\varepsilon)\gamma_{+}^{(i)})^{\ell}\sum_{\ell^{\prime}\geq 0}A_{\ell,\ell^{\prime}}^{(i\to j)}((1+\varepsilon)\gamma_{+}^{(j)})^{\ell^{\prime}}.$ (6.5)
The above sum is a power series that has been introduced in [BBG12b]. More specifically, we deduce from equations (3.16)–(3.18) in [BBG12b] that
$\sum_{\ell\geq 1}\sum_{\ell^{\prime}\geq 0}A_{\ell,\ell^{\prime}}^{(i\to j)}r^{\ell}t^{\ell^{\prime}}=\frac{x_{i}r}{1-x_{i}r-x_{j}t}.$ (6.6)
We are rather interested in the sum over $\ell^{\prime}\geq 0$ when $\ell$ is fixed and with $t=\gamma_{+}^{(j)}$, that is the $r^{k}$-coefficient in the bivariate power series. By expanding in $r$ the ratio on the right-hand side of (6.6) at $t=(1+\varepsilon)\gamma_{+}^{(j)}$, we get that
$\sum_{\ell^{\prime}\geq 0}A_{\ell,\ell^{\prime}}^{(i\to j)}((1+\varepsilon)\gamma_{+}^{(j)})^{\ell^{\prime}}=\bigg(\frac{x_{i}}{1-(1+\varepsilon)x_{j}\gamma_{+}^{(j)}}\bigg)^{\ell+1}.$
Note that the above expansions make sense for $\varepsilon>0$ small enough since $x_{j}\gamma_{+}^{(j)}\leq 1/2$ by 6.1. Thus, from (6.5) we end up with
$K_{\ell}^{(i)}\leq C\bigg(\frac{(1+\varepsilon)x_{i}\gamma_{+}^{(i)}}{1-(1+\varepsilon)x_{j}\gamma_{+}^{(j)}}\bigg)^{\ell}.$ (6.7)
By 6.1, we can choose $\varepsilon>0$ small enough so that the above ratio is in $(0,1)$, which concludes the proof of 1.4 in the case of $\mathcal{S}_{q}$.
Polynomial decay at the self-dual point $(s,p_{0})=(x_{c},p_{\mathrm{sd}}(q))$.
We now complete the proof of 1.4 by showing our last claim about the self-dual regime. We start again from (6.4) which, at the self-dual point, becomes
$K_{\ell}=nF_{\ell}\sum_{\ell^{\prime}\geq 0}A_{\ell,\ell^{\prime}}F_{\ell^{\prime}}.$ (6.8)
But, at the self-dual point, the partition function is explicitly given by our 1.1 as
$F_{k}:=\int_{\gamma_{-}}^{\gamma_{+}}\rho(y)y^{k}\mathrm{d}y,\quad k\geq 1,$
where $(\gamma_{-},\gamma_{+},\rho)$ are also explicit. We may then plug this expression into (6.8) and swap the sum and the integral to get
$K_{\ell}=nF_{\ell}\int_{\gamma_{-}}^{\gamma_{+}}\mathrm{d}y\cdot\rho(y)\sum_{\ell^{\prime}\geq 0}A_{\ell,\ell^{\prime}}y^{\ell^{\prime}}.$
Again we can use the general formula (6.6) to reduce the expression to
$K_{\ell}=nF_{\ell}\int_{\gamma_{-}}^{\gamma_{+}}\rho(y)\Big(\frac{x_{c}}{1-x_{c}y}\Big)^{\ell}\mathrm{d}y.$
Now recall from 3.6 that $x_{c}=1/\gamma_{+}$, so that
$K_{\ell}=nF_{\ell}\gamma_{+}^{-\ell}\int_{\gamma_{-}}^{\gamma_{+}}\rho(y)(2-y/\gamma_{+})^{-\ell}\mathrm{d}y.$
We now use the change of variables $u:=\ell(\gamma_{+}-y)$ in the above integral, which yields
$K_{\ell}=nF_{\ell}\gamma_{+}^{-\ell}\int_{0}^{(\gamma_{+}-\gamma_{-})\ell}\rho(\gamma_{+}-u/\ell)\Big(1+\frac{u}{\gamma_{+}\ell}\Big)^{-\ell}\frac{\mathrm{d}u}{\ell}.$
Finally, using the expression for $\rho$ in 1.1 and dominated convergence, we see that there exists a constant $C>0$ such that
$K_{\ell}\sim CF_{\ell}\frac{\gamma_{+}^{-\ell}}{\ell^{2-\theta}}\quad\text{as }\ell\to\infty.$
We conclude from 1.2 that $K_{\ell}$ has the polynomial decay claimed in 1.4. ∎
@@ Appendix A Computation of inverse Fourier transform
Such integrals can be explicitly computed in terms of Beta functions: indeed, if $a,b\in\mathbb{C}$ with $\Re(a)>|\Re(b)|$ then (see [20, (5.12.7)])